### v2

# Whole-body Motion Planning and Safety-Critical Control for Aerial Manipulation

Abstract
Aerial manipulation combines the maneuverability of multirotors with the dexterity of robotic arms to perform complex tasks in cluttered spaces. Yet planning safe, dynamically feasible trajectories remains difficult due to whole-body collision avoidance and the conservativeness of common geometric abstractions such as bounding boxes or ellipsoids. We present a whole-body motion planning and safety-critical control framework for aerial manipulators built on superquadrics (SQs). Using an SQ-plus-proxy representation, we model both the vehicle and obstacles with differentiable, geometry-accurate surfaces. Leveraging this representation, we introduce a maximum-clearance planner that fuses Voronoi diagrams with an equilibrium-manifold formulation to generate smooth, collision-aware trajectories. We further design a safety-critical controller that jointly enforces thrust limits and collision avoidance via high-order control barrier functions. In simulation, our approach outperforms sampling-based planners in cluttered environments, producing faster, safer, and smoother trajectories and exceeding ellipsoid-based baselines in geometric fidelity. Actual experiments on a physical aerial-manipulation platform confirm feasibility and robustness, demonstrating consistent performance across simulation and hardware settings. The video can be found at https://youtu.be/hQYKwrWf1Ak.
keywords:
Aerial manipulators, whole-body motion planning, collision avoidance, cluttered environments, Voronoi diagrams, safety-critical control.
thanks: Corresponding authorthanks: This work was partly supported by the Institute of Information & Communications Technology Planning & Evaluation(IITP)-ITRC(Information Technology Research Center) grant funded by the Korea government(MSIT)(IITP-2025-RS-2024-00437268, 50%) and the National Research Foundation, Singapore, under the NRF Medium Sized Centre scheme (CARTIN, 50%).
1 Introduction
Aerial manipulation combines the maneuverability of unmanned aerial vehicles with the dexterity of robot arms. Leveraging this synergy, numerous studies have been conducted on diverse applications, including door opening (lee2020aerial), plug-pulling (byun2021stability), non-destructive testing (NDT) (d2025semi), painting (susbielle2025pointillism), and pick-and-place (blochlinger2025reliable).
Bringing such applications into real-world operation requires prioritizing safety –most notably, preventing unintended collisions. This is especially important for aerial manipulators, whose end effectors can reach targets in tight or occluded spaces that would be inaccessible without arm actuation, thereby increasing both opportunity and collision risk. as shown in Fig. 1. Furthermore, in the control level, both dynamic feasibility and instantaneous collision avoidance must be satisfied, as the aerial manipulator may be unable to follow the desired trajectory generated from planners due to the limitations in control performances. However, very few studies have simultaneously addressed collision avoidance between the robot arm’s linkages and surrounding obstacles and the system’s dynamic feasibility.
Figure 1: Robot arm’s actuation for collision avoidance in a cluttered environment.
1.1 Related Works
There have been several studies on collision avoidance of aerial manipulators. In seo2017locally, the authors present a locally optimal trajectory planning method for aerial manipulation in constrained environments, using a signed distance metric and a convex mesh representation to model obstacles. In cao2024motion, the authors propose a motion planning approach for aerial pick-and-place tasks that incorporates collision avoidance constraints between the movable range of the robot arm and surrounding obstacles. Additionally, zhang2025end introduces an end-effector-oriented motion planner for aerial manipulators, which accounts for system dynamics and collision avoidance within constrained spaces. However, these methods only focus on the poses of the multirotor and the end-effector, not the collisions between the robot arm’s linkages and obstacles.
To prevent possible collisions between the robot arm’s linkages and the surrounding environment, there exist other studies that investigate collision avoidance methods that explicitly consider the interaction between these linkages and obstacles. liu2024coordinated presents a model predictive control framework for safe and compliant aerial manipulation, employing simple box constraints on the states of both the multirotor and the robot arm. Additionally, lee2025autonomous proposes a whole-body planning method for omnidirectional aerial manipulators, where both the multirotor platform and the robot arm’s linkages are represented as ellipsoids. Despite these advances, such approaches rely on bounding box or ellipsoid-based geometric representations for aerial manipulators and obstacles, which increase conservativeness when navigating through narrow passages.
To overcome the conservative issue, several studies have developed sampling-based methods for collision avoidance in aerial manipulation. In kim2019sampling, the authors present a rapidly-exploring random trees star (RRT*) -based planner for collision-free aerial pick-and-place tasks, and yavari2022interleaved introduces an RRT*-based approach that divides operational modes into approach and manipulation phases. However, sampling-based methods are generally more time-consuming than approaches based on geometric representations.
1.2 Contributions
In this paper, we propose a whole-body motion planning and a safety-critical control of an aerial manipulator using superquadrics and proxies, wherein both the multirotor platform and the robot arm’s linkages independently achieve collision avoidance with surrounding obstacles. To this end, we first derive the dynamic model of the aerial manipulator attaching a three degree-of-freedom (DOF) robot arm on a fully actuated hexarotor. Then, we design a maximum-clearance whole-body motion planner by integrating the SQ–proxy representation with Voronoi diagrams on an equilibrium manifold (campolo2025geometric). Building upon this planner, a controller complying with propeller thrust limits and collision avoidance is constructed, where the controller framework includes a disturbance observer (DOB)-based control law presented in back2009inner. In particular, our main contributions are arranged as follows:
  • To the best of the authors’ knowledge, this work is the first to employ a superquadric–proxy representation for aerial manipulators, enabling accurate yet efficient whole-body modeling and differentiable distance evaluation for collision avoidance.
  • We design a safety-critical framework that integrates the SQ–proxy representation with Voronoi-based maximum-clearance planning and control barrier functions (CBFs), ensuring dynamic feasibility and instantaneous collision avoidance under thrust limits.
  • Comparative simulations and real-world experiments demonstrate that our approach achieves faster, safer, and smoother trajectories than sampling-based planners, while surpassing ellipsoid-based baselines in geometry accuracy.
1.3 Outline
This paper is outlined as follows. In Section 2, we formulate the dynamic model of our fully actuated aerial manipulator, and the path planning method based on superquadrics formulation, Voronoi diagram, and potential function is introduced in Section 3. In Section 4, we present a safety-critical controller guaranteeing thrust limits and instantaneous collision avoidance. Section 5 covers the comparative simulations of the proposed method, and our method is also validated in a real-world experiment with a fully actuated aerial manipulator in Section 6.
1.4 Notations
[…]
2 Aerial Manipulator System
As a platform, we utilize an aerial manipulator configured with a fully actuated multirotor with six tilted motors and a 3-DOF robot arm, as illustrated in Fig. 2. We let $\mathscr{F}_{W}$ and $\mathscr{F}_{b}$ denote two key coordinate frames, the Earth-fixed and multirotor body frames, respectively, then the position and ZYX Euler angles of $\mathscr{F}_{b}$ w.r.t. $\mathscr{F}_{W}$ are defined as $\boldsymbol{p}$ and $\boldsymbol{\phi}\in{\mathbb{R}}^{3}$, respectively. Since we consider the movement of the robot arm as a part of external disturbances, the generalized coordinate of the aerial manipulator, $\boldsymbol{q}\in{\mathbb{R}}^{6}$, is defined as $\boldsymbol{p};\boldsymbol{\phi}]$.
Figure 2: An aerial manipulator configured with a fully actuated multirotor and a 3-DOF robot arm with the thrust vector, $\boldsymbol{T}\in{\mathbb{R}}^{6}$, the joint angles of the robot arm, $\theta_{1};\theta_{2};\theta_{3}]$, and the Earth-fixed and multirotor frames, $\mathscr{F}_{W}$ and $\mathscr{F}_{b}$.
According to byun2025safety, the Euler-Lagrange model of the aerial manipulator is formulated as follows:
$\begin{split}\boldsymbol{M}(\boldsymbol{\phi})\boldsymbol{\ddot{q}}+\boldsymbol{C}(\boldsymbol{\phi},\boldsymbol{\dot{\phi}})+\boldsymbol{G}=\boldsymbol{\tau}+\boldsymbol{\tau_{ext}}\end{split}$ (1)
where $\boldsymbol{\tau}$ and $\boldsymbol{\tau_{ext}}\in{\mathbb{R}}^{6}$ represent the generalized control wrench and external disturbances, respectively, and
$\begin{split}\boldsymbol{M}(\boldsymbol{\phi})\triangleq&\begin{bmatrix}m\boldsymbol{I_{3}}&\boldsymbol{0_{3\times 3}}\\ \boldsymbol{0_{3\times 3}}&\boldsymbol{Q^{\top}}\boldsymbol{J}\boldsymbol{Q}\end{bmatrix},\\ \boldsymbol{C}(\boldsymbol{\phi},\boldsymbol{\dot{\phi}})\triangleq&\begin{bmatrix}\boldsymbol{0_{3\times 1}}\\ \boldsymbol{Q^{\top}}(\boldsymbol{J}\boldsymbol{\dot{Q}}\boldsymbol{\dot{\phi}}+[\boldsymbol{Q}\boldsymbol{\dot{\phi}}]_{\times}\boldsymbol{J}\boldsymbol{Q}\boldsymbol{\dot{\phi}})\end{bmatrix},\ \boldsymbol{G}\triangleq\begin{bmatrix}mg\boldsymbol{e_{3}}\\ \boldsymbol{0_{3\times 1}}\end{bmatrix}\end{split}$
with the mass and moment of inertia of the aerial manipulator, $m$ and $\boldsymbol{J}\in{\mathbb{R}}^{3\times 3}$, and the gravitational acceleration, $g$. Also, we let $\boldsymbol{Q}\in{\mathbb{R}}^{3\times 3}$ denote the mapping matrix satisfying $\boldsymbol{\omega}=\boldsymbol{Q}\boldsymbol{\dot{\phi}}$ where $\boldsymbol{\omega}\in{\mathbb{R}}^{3}$ represents the angular velocity of the vehicle w.r.t. $\mathscr{F}_{W}$ expressed in $\mathscr{F}_{b}$. Meanwhile, $\boldsymbol{\tau}$ and $\boldsymbol{T}\triangleq[T_{1};\cdots;T_{6}]$ satisfy:
$\begin{split}\boldsymbol{\tau}=\boldsymbol{B}(\boldsymbol{\phi})\boldsymbol{T}\end{split}$ (2)
where
$\begin{split}&\boldsymbol{B}(\boldsymbol{\phi})\triangleq\textrm{blkdiag}\{\boldsymbol{R},\boldsymbol{Q^{\top}}\}\\ &\qquad\times\begin{bmatrix}\tfrac{1}{2}s\alpha_{p}&-s\alpha_{p}&\tfrac{1}{2}s\alpha_{p}&\tfrac{1}{2}s\alpha_{p}&-s\alpha_{p}&\tfrac{1}{2}s\alpha_{p}\\ -\tfrac{\sqrt{3}}{2}s\alpha_{p}&0&\tfrac{\sqrt{3}}{2}s\alpha_{p}&-\tfrac{\sqrt{3}}{2}s\alpha_{p}&0&\tfrac{\sqrt{3}}{2}s\alpha_{p}\\ c\alpha_{p}&c\alpha_{p}&c\alpha_{p}&c\alpha_{p}&c\alpha_{p}&c\alpha_{p}\\ -\tfrac{1}{2}P_{1}&-P_{1}&-\tfrac{1}{2}P_{1}&\tfrac{1}{2}P_{1}&P_{1}&\tfrac{1}{2}P_{1}\\ \tfrac{\sqrt{3}}{2}P_{1}&0&-\tfrac{\sqrt{3}}{2}P_{1}&-\tfrac{\sqrt{3}}{2}P_{1}&0&\tfrac{\sqrt{3}}{2}P_{1}\\ P_{2}&-P_{2}&P_{2}&-P_{2}&P_{2}&-P_{2}\end{bmatrix},\\ &P_{1}\triangleq Lc\alpha_{p}+k_{f}s\alpha_{p},\ P_{2}\triangleq Ls\alpha_{p}-k_{f}c\alpha_{p}\end{split}$
where $L$, $\alpha_{p}$ and $k_{f}$ represent the length from the vehicle’s origin to each propeller, fixed tilt angle of each motor and thrust-to-torque coefficient, respectively.
To address the effect of model uncertainties, (1) is rearranged as:
$\begin{split}\boldsymbol{\hat{M}}(\boldsymbol{\phi})\boldsymbol{\ddot{q}}+\boldsymbol{\hat{C}}(\boldsymbol{\phi},\boldsymbol{\dot{\phi}})+\boldsymbol{\hat{G}}=\boldsymbol{\tau}+\boldsymbol{d}\end{split}$ (3)
where $\boldsymbol{\hat{M}}(\boldsymbol{\phi})$, $\boldsymbol{\hat{C}}(\boldsymbol{\phi},\boldsymbol{\dot{\phi}})$ and $\boldsymbol{\hat{G}}$ are the nominal values of $\boldsymbol{M}(\boldsymbol{\phi})$, $\boldsymbol{C}(\boldsymbol{\phi},\boldsymbol{\dot{\phi}})$ and $\boldsymbol{G}$, respectively, and the lumped disturbance $\boldsymbol{d}\in{\mathbb{R}}^{6}$ is derived as follows:
$\boldsymbol{d}\triangleq(\boldsymbol{\hat{M}}(\boldsymbol{\phi})-\boldsymbol{M}(\boldsymbol{\phi}))\boldsymbol{\ddot{q}}\\ +\boldsymbol{\hat{C}}(\boldsymbol{\phi},\boldsymbol{\dot{\phi}})-\boldsymbol{C}(\boldsymbol{\phi},\boldsymbol{\dot{\phi}})+\boldsymbol{\hat{G}}-\boldsymbol{G}+\boldsymbol{\tau_{ext}}.$
3 Whole-body motion planning
Building upon yang2024path; yang2025planning and kana2021human, we model the entire aerial manipulator system as well as environmental obstacles using SQs. The interactions between any two SQs are represented by proxies. Meanwhile, the end-effector of the robot arm follows Voronoi edges, while collision avoidance is also handled via proxies. The entire pipeline is illustrated in Fig. 3.
Figure 3: Overall diagram of planning and control of aerial manipulator.
3.1 Superquadrics and proxies
SQs offer a computationally efficient representation of real-world object geometry (jaklic2000segmentation). We define an SQ as follows:
$\begin{split}F_{3d}(x,y,z)=\Big((\tfrac{x}{a_{1}})^{\tfrac{2}{\epsilon_{2}}}+(\tfrac{y}{a_{2}})^{\tfrac{2}{\epsilon_{2}}}\Big)^{\tfrac{\epsilon_{2}}{\epsilon_{1}}}+(\tfrac{z}{a_{3}})^{\tfrac{2}{\epsilon_{1}}}-1\end{split}$ (4)
where $\epsilon_{1}$ and $\epsilon_{2}$ represent the exponents of the 3D superellipse controlling its shape, while $a_{1},a_{2}$ and $a_{3}$ are size-controlling parameters. Also, we introduce the concept of a proxy sliding along the SQ boundary, parameterized by the angular variable $\boldsymbol{\gamma_{3D}}$:
${\boldsymbol{\mathbf{p}}}(\boldsymbol{\gamma_{3D}})=\begin{bmatrix}a_{1}\cos^{\epsilon_{1}}\gamma_{3D,1}\cos^{\epsilon_{2}}\gamma_{3D,2}\\ a_{1}\cos^{\epsilon_{1}}\gamma_{3D,1}\sin^{\epsilon_{2}}\gamma_{3D,2}\\ a_{3}\sin^{\epsilon_{1}}\gamma_{3D,1}\end{bmatrix}$,
$\boldsymbol{\gamma_{3D}}\in[-\tfrac{\pi}{2},\ \tfrac{\pi}{2}]\times[-\pi,\pi]$ (5)
where $\cos^{\ast}(\cdot)=\textrm{sign}(\cos(\cdot))|\cos(\cdot)|^{\ast}$ and $\sin^{\ast}(\cdot)=\textrm{sign}(\sin(\cdot))|\sin(\cdot)|^{\ast}$.
To ensure the safety and stability of the multirotor, excessive tilting is restricted. Under this assumption, the task can be modeled in a 2D scenario. Accordingly, Eqs. 4 and 3.1 reduce to their 2D formulations, involving only the parameters $a_{1},a_{2},\epsilon,$ and $\gamma\in[-\pi,\pi]$, as presented in yang2025planning.
3.2 Maximum clearance path planning
To enhance the safety of the aerial manipulator system, we adopt the Voronoi diagram as the foundation of our motion planner, since it naturally provides maximal clearance from obstacles in arslan2019sensor. Following the approach in yang2024path, the Voronoi diagram is constructed using SQs and proxies.
3.2.1 (1) Voronoi diagram
A Voronoi diagram partitions the Euclidean space $E\subset\mathbb{R}^{n}$ into $m$ regions $\{V_{1},V_{2},\ldots,V_{m}\}$, commonly referred to as cells. Each cell is defined such that every point within it is closest to the corresponding disjoint convex obstacle SQ (i.e., $d(SQ_{i},SQ_{j})>0,\ \forall i\neq j$). Based on the maximum-margin separating hyperplane, the construction is given by
$HP_{i,j}$ $:=\left\{{\boldsymbol{\mathbf{p}}}\in E\ \middle|\ \|{\boldsymbol{\mathbf{p}}}-{\boldsymbol{\mathbf{p}}}(\gamma_{i})\|=\|{\boldsymbol{\mathbf{p}}}-{\boldsymbol{\mathbf{p}}}(\gamma_{j})\|\right\},$ (6a)
$V_{i}$ $:=\left\{{\boldsymbol{\mathbf{p}}}\in W\middle|\ \|{\boldsymbol{\mathbf{p}}}-{\boldsymbol{\mathbf{p}}}(\gamma_{i})\|\leq\|{\boldsymbol{\mathbf{p}}}-{\boldsymbol{\mathbf{p}}}(\gamma_{j})\|\right\},$ (6b)
where $HP_{i,j}$ denotes the hyperplane (or boundary) shared by two Voronoi regions $V_{i}$ and $V_{j}$, computed from Eq. 3.1. Traveling along $HP_{i,j}$ guarantees that the path remains equidistant from the obstacles associated with $V_{i}$ and $V_{j}$, thus forming a so-called maximum-clearance path.
3.2.2 (2) Global planner
We first define a bounding box (world space $W$) to appropriately represent the workspace. Consequently, each region $V_{i}$ is expressed as the intersection of all hyperplanes defining its polyhedral cell within the bounding box:
$V_{i}=\bigcup_{j=0}^{m}W\cap HP_{i,j}.$ (7)
Once all hyperplanes are computed, the Voronoi diagram is represented as a graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$, where each $\mathcal{V}_{i}$ corresponds to a polygon defined by multiple vertices and edges. Each vertex is represented as a node $v_{i}\in\mathcal{V}$, and nodes are connected by edges in $\mathcal{E}$. The weight assigned to each edge corresponds to the Euclidean distance between its two nodes. Finally, classical graph search algorithms can be applied to compute a solution path, denoted by $(\mathcal{V}_{\mathrm{sol}},\mathcal{E}_{\mathrm{sol}})$, consisting of selected nodes, edges, and their orientations.
3.3 Planning on equilibrium manifold
Following yang2025planning; campolo2025geometric, we formulate the planning task as a manipulation potential function $W:\mathcal{Z}\times\mathcal{U}\rightarrow\mathbb{R}$, where $\boldsymbol{\mathbf{u}}$} denotes the attracting pose of the robotic end effector in control space $\mathcal{U}$, and $\boldsymbol{\mathbf{z}}$} denotes the state of the aerial manipulator system in state space $\mathcal{Z}$. The objective is to generate a smooth trajectory such that the end effector reaches the target while the entire system avoids collisions. This problem can be interpreted as finding an equilibrium state $\boldsymbol{\mathbf{z}}$$^{*}$ on the equilibrium manifold (EM) of the system:
$\mathcal{M}^{eq}:=\{({\boldsymbol{\mathbf{z}}},{\boldsymbol{\mathbf{u}}})\in\mathcal{Z}\times\mathcal{U}|\partial_{{\boldsymbol{\mathbf{z}}}}W({\boldsymbol{\mathbf{z}}},{\boldsymbol{\mathbf{u}}})={\boldsymbol{\mathbf{0}}}\}$ (8)
We define $\partial_{{\boldsymbol{\mathbf{q}}}}W\equiv[\partial_{q_{1}}W,\ldots,\partial_{q_{a}}W]^{T}$, where the partial operator is defined as $\partial_{{\boldsymbol{\mathbf{q}}}}=[\partial_{q_{1}},\ldots,\partial_{q_{a}}]^{T}$. Since the equilibrium state $\boldsymbol{\mathbf{z}}$$^{*}$ is implicitly defined, we adopt the adaptive ODE introduced in yang2025planning to update $\boldsymbol{\mathbf{z}}$$^{*}$, where $\eta$ represents step size in the ODE :
$\dot{{\boldsymbol{\mathbf{z}}}}=-(\partial^{2}_{\boldsymbol{\mathbf{zz}}}W)^{-1}\partial^{2}_{\boldsymbol{\mathbf{uz}}}W\dot{{\boldsymbol{\mathbf{u}}}}-\eta(\partial^{2}_{\boldsymbol{\mathbf{zz}}}W)^{-1}\partial_{\boldsymbol{\mathbf{z}}}W$ (9)
3.4 System modeling
Figure 4: System representation using SQs and proxies. The end effector of the robot arm aligns with the Voronoi edge to ensure collision-free motion.
As illustrated in Fig. 4, the aerial manipulator operates in a 2D space. Define ${\boldsymbol{\mathbf{z}}}_{mr}\in SE(2)$, and consider a two-link robot arm with two motors rotating along the $z$-axis with angles $\theta_{1}$ and $\theta_{3}$. The multirotor with six blades is modeled by six SQs, while the two links are modeled by two SQs, resulting in a total of $SQ_{i},\ i=1,\ldots,8$. Each SQ $SQ_{i}$ is associated with a proxy pair $\gamma_{i},\gamma_{j}$ corresponding to obstacle $SQ_{j}$. Thus, the state variables of the entire system are denoted as ${\boldsymbol{\mathbf{z}}}=[{\boldsymbol{\mathbf{z}}}_{sys},\Gamma]$, where ${\boldsymbol{\mathbf{z}}}_{sys}=[{\boldsymbol{\mathbf{z}}}_{mr},\theta_{1},\theta_{3}]$ represents the combined states of the multirotor and joint, and $\Gamma=[\ldots,\gamma_{i},\ldots,\gamma_{j},\ldots]$ denotes all proxies. The pose of the end effector ${\boldsymbol{\mathbf{z}}}_{EEF}\in SE(2)$ is obtained from forward kinematics.
3.4.1 (1) collision avoidance
The proxy-based potential is used for collision avoidance and proxy dynamics, and is defined as
$W_{proxy}=\sum_{i}\sum_{j}\frac{1}{2}k(d_{ij}-d^{\prime})\|{\boldsymbol{\mathbf{p}}}(\gamma_{i})-{\boldsymbol{\mathbf{p}}}(\gamma_{j})\|^{2}$ (10)
where $d_{ij}=F\!\left({}^{j}p(\gamma_{i})\right)$ represents the evaluation of $p(\gamma_{i})$ in the frame of obstacle $SQ_{j}$, serving as the 2D version of the inside–outside function in Eq. 4. Here, $k(d)$ is a nonlinear stiffness function: it increases as the multirotor or arm approaches an obstacle and decreases when farther away. The parameter $d^{\prime}$ specifies the safety margin for obstacle avoidance:
$k(d)=k_{\text{min}}+\frac{1-\tanh(d/d_{0})}{2}k_{\text{max}}.$ (11)
3.4.2 (2) Target following
To ensure that the end effector achieves the desired pose, we define an attracting frame ${\boldsymbol{\mathbf{u}}}=[u_{x},u_{y},u_{\theta}]\in SE(2)$, where $u_{x}$ and $u_{y}$ follow $\mathcal{E}_{sol}$, and $u_{\theta}$ aligns with the normal of the Voronoi edge. A virtual spring is applied to connect $\boldsymbol{\mathbf{u}}$} with the end effector ${\boldsymbol{\mathbf{z}}}_{EEF}$. The attracting potential is given by
$W_{tgt}=\frac{1}{2}({\boldsymbol{\mathbf{u}}}-{\boldsymbol{\mathbf{z}}}_{EEF})^{T}{\boldsymbol{\mathbf{K}}}_{tgt}({\boldsymbol{\mathbf{u}}}-{\boldsymbol{\mathbf{z}}}_{EEF})$ (12)
Finally, the manipulation potential is $W=W_{proxy}+W_{tgt}$.
3.5 Local planner
The global planner, based on the Voronoi diagram, returns a set of discrete nodes and edges $(\mathcal{V}_{sol},\mathcal{E}_{sol})$. Each via point on this path is represented as a discrete attraction pose ${\boldsymbol{\mathbf{u}}}_{i}$. To effectively guide the aerial manipulator system, we require a continuous and smooth reference trajectory ${\boldsymbol{\mathbf{z}}}^{*}(s),\ s\in[0,1]$. To obtain such a trajectory, we integrate the following augmented ODE system from $s=0$ to $s=1$:
$\tfrac{\partial\boldsymbol{{\boldsymbol{\mathbf{u}}}}}{\partial s}$ $={\boldsymbol{\mathbf{u}}}_{i+1}-{\boldsymbol{\mathbf{u}}}_{i},$ (13a)
$\tfrac{\partial{\boldsymbol{\mathbf{z}}}_{sys}}{\partial s}$ $=-(\partial^{2}_{\boldsymbol{\mathbf{zz}}}W)^{-1}\partial^{2}_{\boldsymbol{\mathbf{uz}}}W\tfrac{\partial{\boldsymbol{\mathbf{z}}}_{sys}}{\partial s}-\eta(\partial^{2}_{\boldsymbol{\mathbf{zz}}}W)^{-1}\partial_{\boldsymbol{\mathbf{z}}}W,$ (13b)
$\tfrac{\partial{\boldsymbol{\mathbf{\Gamma}}}}{\partial s}$ $=-\alpha\partial_{{\boldsymbol{\mathbf{\gamma}}}}{W},$ (13c)
where initial condition comes from the aerial manipulator’s current state.
3.6 Target Pose Calculation
From the results of the local planner, the target pose of the multirotor, $\boldsymbol{q_{t}}$, and the target angles of the robot arm’s servo motors, $\boldsymbol{\theta_{t}}$, are calculated as follows:
$\begin{split}\boldsymbol{q_{t}}=&[z_{sys,1}(s(t));z_{sys,2}(s(t));h_{t};0;0;z_{sys,3}(s(t))]\\ \boldsymbol{\theta_{t}}=&[z_{sys,4}(s(t));0;z_{sys,5}(s(t))]\end{split}$ (14)
where the time allocated path parameter $s(t)$ is calculated as $\tfrac{1}{T_{d}}t$ with the desired task execution time $T_{d}>0$, and $h_{t}$ represents the target altitude of the multirotor.
4 Safety-Critical Controller
This section presents a safety-critical controller that tracks the desired trajectories of both the multirotor and the robot arm, subject to two safety constraints. The first constraint enforces propeller-thrust limits: when $\boldsymbol{q_{t}}$ is substantially far from the current pose $\boldsymbol{q}$, the required control effort can drive one or more rotor thrust beyond their physical bounds. The second constraint enforces instantaneous collision avoidance. This addresses cases where tracking errors between $(\boldsymbol{q},\boldsymbol{\theta})$ and $(\boldsymbol{q_{t}},\boldsymbol{\theta_{t}})$ arises due to limited control authority, even when the planned global path is collision-free. As depicted in Fig. 3, the proposed controller is divided into two sequences: outer- and inner-loop control laws.
4.1 Inner-Loop Control Law
To cope with three performance deteriorating factors, i.e., movement of the attached robot arm, model uncertainties, and time-varying external disturbance, we adopt a DOB-based controller extending the control law introduced in kim2017robust to incorporate 6 DOF pose.
Let $\boldsymbol{q_{d}}$ and $\boldsymbol{\theta_{d}}$ denote the desired pose of the multirotor and the desired joint angles of the robot arm, respectively, which the multirotor and robot arm attempt to track. However, these may differ from the target values, $\boldsymbol{q_{t}}$ and $\boldsymbol{\theta_{t}}$, since an outer-loop controller is also designed to adjust the target values in accordance with the safety constraints. This controller will be introduced in a later subsection.
With $\boldsymbol{q_{d}}$ and $\boldsymbol{\theta_{d}}$, our inner-loop control law is formulated as follows:
$\begin{split}\boldsymbol{T}=&\boldsymbol{B^{-1}}(\boldsymbol{\phi})\{\boldsymbol{\hat{M}}(\boldsymbol{\phi})(\boldsymbol{K_{d}}\boldsymbol{\dot{e}}+\boldsymbol{K_{p}}\boldsymbol{e})\\ &\qquad\qquad\qquad+\boldsymbol{\hat{C}}(\boldsymbol{\phi},\boldsymbol{\dot{\phi}})+\boldsymbol{\hat{G}}-\boldsymbol{\hat{d}}\}\\ \boldsymbol{\hat{d}}=&-\boldsymbol{\hat{M}}(\boldsymbol{\phi})(\boldsymbol{E_{1}}\boldsymbol{p_{dob}}-\boldsymbol{E_{2}}\boldsymbol{\dot{q}_{dob}})+\boldsymbol{\hat{C}}(\boldsymbol{\phi},\boldsymbol{\dot{\phi}})+\boldsymbol{\hat{G}}\\ \boldsymbol{\dot{q}_{dob}}=&\boldsymbol{A_{dob}}\boldsymbol{q_{dob}}+\boldsymbol{B_{dob}}\boldsymbol{q}\\ \boldsymbol{\dot{p}_{dob}}=&\boldsymbol{A_{dob}}\boldsymbol{p_{dob}}+\boldsymbol{B_{dob}}\boldsymbol{M^{-1}(\boldsymbol{\phi})}\boldsymbol{B}(\boldsymbol{\phi})\boldsymbol{T}\end{split}$ (15)
where $\boldsymbol{e}\triangleq\boldsymbol{q_{d}}-\boldsymbol{q}$, $\boldsymbol{E_{1}}\triangleq\begin{bmatrix}\boldsymbol{I_{6}}&\boldsymbol{0_{6\times 6}}\end{bmatrix}$ and $\boldsymbol{E_{2}}\triangleq\begin{bmatrix}\boldsymbol{0_{6\times 6}}&\boldsymbol{I_{6}}\end{bmatrix}$ with user-defined positive definite matrices $\boldsymbol{K_{p}},\boldsymbol{K_{d}}\in\mathbb{R}^{6\times 6}$. Also, $\boldsymbol{\hat{d}}$ represents the estimated lumped disturbance, and
$\begin{split}\boldsymbol{A_{dob}}\triangleq&\begin{bmatrix}\boldsymbol{0_{6\times 6}}&\boldsymbol{I_{6}}\\ -\boldsymbol{\varepsilon_{dob}^{-2}}\boldsymbol{a_{0}}&-\boldsymbol{\varepsilon_{dob}^{-1}}\boldsymbol{a_{1}}\end{bmatrix},\quad\boldsymbol{B_{dob}}\triangleq\begin{bmatrix}\boldsymbol{0_{6\times 6}}\\ \boldsymbol{\varepsilon_{dob}^{-2}}\boldsymbol{a_{0}}\end{bmatrix}\end{split}$
with user-defined matrices $\boldsymbol{a_{0}}\triangleq\textrm{diag}\{a_{0,1},\cdots,a_{0,6}\}$, $\boldsymbol{a_{1}}\triangleq\textrm{diag}\{a_{1,1},\cdots,a_{1,6}\}$ and $\boldsymbol{\varepsilon_{dob}}\triangleq\textrm{diag}\{\varepsilon_{dob,1},\cdots,\varepsilon_{dob,6}\}$ under $0<a_{0,i},a_{1,i}$, $\tfrac{1}{2}<\tfrac{a_{0,i}}{a^{2}_{1,i}}$ and $0<\varepsilon_{dob,i}<1$ $\forall i=1,\cdots,6$.
4.2 Safety Constraints
The two safety conditions are formulated as follows:
4.2.1 (1) Thrust Limit
The thrust values have their lower and upper limits, $\underline{T}$ and $\bar{T}$, as follows:
$\begin{split}\underline{T}\boldsymbol{1_{6}}\leq\boldsymbol{T}\leq\bar{T}\boldsymbol{1_{6}}\end{split}$ (16)
where $\boldsymbol{1_{6}}\triangleq[1\ 1\ 1\ 1\ 1\ 1]^{\top}$. If one of the propellers reaches its limit, the multirotor may lose controllability.
4.2.2 (2) Collision Avoidance
To compensate for possible target-tracking errors in the proposed inner-loop controller, an additional strategy at the control level is necessary for collision avoidance, depending on the aerial manipulator’s pose and the robot arm’s joint angles in real time.
To derive a mathematical expression for this constraint, we utilize superquadrics as defined in Eq.4. For $SQ_{i}$ (one of the vehicle SQs) and $SQ_{j}$ (one of the obstacle SQs) ($i=1.\cdots,8$, $j=1.\cdots,m$), the closest points follows Eq. 3.1. We let $\boldsymbol{\Delta X}\in\mathbb{R}^{3}$ denote the displacement from the center of $SQ_{j}$ to the proxy of $SQ_{i}$ as follows:
$\begin{split}&\boldsymbol{\Delta X}\triangleq[\Delta x;\Delta y;\Delta z]\\ &\quad\triangleq\boldsymbol{R^{\top}_{j}}(\boldsymbol{p_{i}}-\boldsymbol{p_{j}})+\boldsymbol{R^{\top}_{j}}\boldsymbol{R_{i}}\boldsymbol{p}(\boldsymbol{\gamma_{3D,i}})\end{split}$ (17)
where $\boldsymbol{R_{i}}$ and $\boldsymbol{R_{j}}\in SO(3)$ represent the rotation matrices from $SQ_{i}$ to $\mathcal{F}_{W}$ and from $SQ_{j}$ to $\mathcal{F}_{W}$, respectively, $\boldsymbol{p_{i}}$ and $\boldsymbol{p_{j}}\in\mathbb{R}^{3}$ mean the displacement from $\mathcal{F}_{W}$ to the centers of $SQ_{i}$ and $SQ_{j}$, respectively, and $\boldsymbol{\gamma_{3D,i}}$ is the angular variable of $SQ_{i}$. Then, a control barrier function (CBF) for collision avoidance is defined as follows:
$\begin{split}h_{co}\triangleq\ln(\Big((\tfrac{\Delta x}{{}^{j}a_{1}})^{\tfrac{2}{{}^{j}\epsilon_{2}}}+(\tfrac{\Delta y}{{}^{j}a_{2}})^{\tfrac{2}{{}^{j}\epsilon_{2}}}\Big)^{\tfrac{{}^{j}\epsilon_{2}}{{}^{j}\epsilon_{1}}}+(\tfrac{\Delta z}{{}^{j}a_{3}})^{\tfrac{2}{{}^{j}\epsilon_{1}}})\end{split}$ (18)
where $^{j}\epsilon_{1}$, $^{j}\epsilon_{2}$, $^{j}a_{1}$, $^{j}a_{2}$ and $^{j}a_{3}$ are the parameters of $SQ_{j}$.
According to Eq. 18, $h_{co}$ becomes positive when the proxy of $SQ_{i}$ is located outside the $SQ_{j}$; namely, two SQs do not collide. Thus, $h_{co}$ can be a CBF for the collision avoidance.
4.3 Outer-Loop Control Law
Given the target pose of the aerial manipulator from Eq. 14, the desired pose and twist, $\boldsymbol{q_{d}};\boldsymbol{\theta_{d}}]$ and $\boldsymbol{\dot{q}_{d}};\boldsymbol{\dot{\theta}_{d}}]$, need to be generated based on the two safety conditions.
For the thrust limits, we rearrange Eq. 16 as follows:
$\begin{split}\boldsymbol{A_{\underline{T}}}\boldsymbol{\dot{q}_{d}}\leq\boldsymbol{b_{\underline{T}}},\ \boldsymbol{A_{\bar{T}}}\boldsymbol{\dot{q}_{d}}\leq\boldsymbol{b_{\bar{T}}},\end{split}$ (19)
where $\boldsymbol{A_{\underline{T}}}\triangleq-\boldsymbol{B}^{-1}(\boldsymbol{\phi})\boldsymbol{\hat{M}}(\boldsymbol{\phi})\boldsymbol{K_{d}}$, $\boldsymbol{A_{\bar{T}}}\triangleq-\boldsymbol{A_{\underline{T}}}$ and
[…]
Meanwhile, since $h_{co}$ is the function of $(\boldsymbol{q},\boldsymbol{\theta})$, we need to differentiate it twice to attain the thrust inputs. Hence, Eq. 18 becomes a high-order CBF (xiao2021high), and the collision avoidance condition is formulated as follows:
$\ddot{h}_{co}+2\alpha_{co}\dot{h}_{co}+\alpha^{2}_{co}h_{co}\geq 0$ (20)
where $\alpha_{co}>0$ is a user-defined parameter. With the assumption that $\boldsymbol{\gamma_{3D,i}}$ and $\boldsymbol{\gamma_{3D,j}}$ slowly change with time, the time-derivatives of $h_{co}$ depend on those of $\boldsymbol{p_{i}}$ and $\boldsymbol{R_{i}}$. Therefore, $\ddot{h}_{co}$ and $\dot{h}_{co}$ can be expressed as follows:
$\begin{split}\dot{h}_{co}=&\tfrac{\partial h_{co}}{\partial\boldsymbol{\Delta X}}\boldsymbol{\dot{(\Delta X)}}\\ \ddot{h}_{co}=&\tfrac{\partial h_{co}}{\partial\boldsymbol{\Delta X}}\boldsymbol{\ddot{(\Delta X)}}+\boldsymbol{\dot{(\Delta X)}^{\top}}\tfrac{\partial^{2}h_{co}}{\partial(\boldsymbol{\Delta X})^{2}}\boldsymbol{\dot{(\Delta X)}}.\end{split}$ (21)
where we can find matrices $\boldsymbol{A_{\Delta X}}\in{\mathbb{R}}^{3\times 9}$ and $\boldsymbol{b_{\Delta X}}\in{\mathbb{R}}^{3}$ such that $\boldsymbol{\ddot{(\Delta X)}}=\boldsymbol{A_{\Delta X}}[\boldsymbol{\ddot{q}};\boldsymbol{\ddot{\theta}}]+\boldsymbol{b_{\Delta X}}$ from Eq. 17. Also, by substituting Eq. 15 for Eq. 3, $\boldsymbol{\ddot{q}}$ is rearranged as follows:
$\begin{split}\boldsymbol{\ddot{q}}=\boldsymbol{K_{d}}\boldsymbol{\dot{q}_{d}}-\boldsymbol{K_{d}}\boldsymbol{\dot{q}}+\boldsymbol{K_{p}}\boldsymbol{e}-\boldsymbol{\hat{M}^{-1}}(\boldsymbol{\phi})\boldsymbol{\tilde{d}}\end{split}$ (22)
where $\boldsymbol{\tilde{d}}\triangleq\boldsymbol{\hat{d}}-\boldsymbol{d}$. Assume that $\boldsymbol{\ddot{\theta}}\approx\boldsymbol{\ddot{\theta}_{d}}$, then (20) is rearranged as follows:
$\boldsymbol{A_{co}}[\boldsymbol{\dot{q}_{d}};\boldsymbol{\ddot{\theta}_{d}}]+d_{co}\leq b_{co}$ (23)
where
$\begin{split}\boldsymbol{A_{co}}\triangleq&-\tfrac{\partial h_{co}}{\partial\boldsymbol{\Delta X}}\boldsymbol{A_{\Delta X}}\begin{bmatrix}\boldsymbol{K_{d}}&\boldsymbol{0_{6\times 3}}\\ \boldsymbol{0_{3\times 6}}&\boldsymbol{I_{3}}\end{bmatrix},\\ d_{co}\triangleq&\tfrac{\partial h_{co}}{\partial\boldsymbol{\Delta X}}\boldsymbol{A_{\Delta X}}\begin{bmatrix}\boldsymbol{\hat{M}^{-1}}(\boldsymbol{\phi})\boldsymbol{\tilde{d}}\\ \boldsymbol{0_{3\times 1}}\end{bmatrix},\\ b_{co}\triangleq&\tfrac{\partial h_{co}}{\partial\boldsymbol{\Delta X}}\boldsymbol{A_{\Delta X}}\begin{bmatrix}-\boldsymbol{K_{d}}\boldsymbol{\dot{q}}+\boldsymbol{K_{p}}\boldsymbol{e}\\ \boldsymbol{0_{3\times 1}}\end{bmatrix}+\tfrac{\partial h_{co}}{\partial\boldsymbol{\Delta X}}\boldsymbol{b_{\Delta X}}+\\ &\boldsymbol{\dot{(\Delta X)}^{\top}}\tfrac{\partial^{2}h_{co}}{\partial(\boldsymbol{\Delta X})^{2}}\boldsymbol{\dot{(\Delta X)}}+2\alpha_{co}\tfrac{\partial h_{co}}{\partial\boldsymbol{\Delta X}}\boldsymbol{\dot{(\Delta X)}}+\alpha^{2}_{co}h_{co}.\end{split}$
As $d_{co}$ contains the disturbance estimation error, a more conservative collision avoidance constraint is expressed as:
$\boldsymbol{A_{co}}[\boldsymbol{\dot{q}_{d}};\boldsymbol{\ddot{\theta}_{d}}]+\sigma_{co}\leq b_{co}$ (24)
where $|d_{co}|\leq\sigma_{co}$.
Finally, a quadratic programming (QP)-based outer-loop controller is constructed as follows:
[…]
where $\boldsymbol{\dot{q}_{d,ref}}\triangleq\boldsymbol{\Gamma_{q}}(\boldsymbol{q_{t}}-\boldsymbol{q_{d}})$ and $\boldsymbol{\ddot{\theta}_{d,ref}}\triangleq-2\boldsymbol{\Gamma_{\theta}}\boldsymbol{\dot{\theta}_{d}}+\boldsymbol{\Gamma^{2}_{\theta}}(\boldsymbol{\theta_{t}}-\boldsymbol{\theta_{d}})$ with weight matrices $\boldsymbol{Q_{\dot{q}}},\boldsymbol{Q_{\ddot{\theta}}}\in\mathbb{R}^{3\times 3}_{>0}$, and target-following coefficients $\boldsymbol{\Gamma_{q}},\boldsymbol{\Gamma_{\theta}}\in\mathbb{R}^{3\times 3}_{>0}$. Then, the desired pose and joint angles are obtained by integration.
5 Simulation Results
We first evaluate the proposed planner in simulation. Two representative scenarios with various initial conditions are tested, and the results are compared against several typical and widely used baseline planners.
5.1 Results from whole-body motion planner
The results from the whole-body motion planner are shown in Fig. 5, where the two scenarios correspond to environments with trees and pillars respectively. These scenarios are designed such that the target can only be reached by exploiting the thin linkages of the robot arm, thereby validating the necessity of whole-body planning.
(a) Tree-like obstacles.
(b) Pillar-like obstacles.
Figure 5: Simulation results from the whole-body motion planner. Blue transparent polygons represent Voronoi regions $V_{i}$ separated by grey hyperplanes $HP_{i,j}$, computed via proxies between obstacles. The aerial manipulator and obstacles are modeled by SQs (black). The red point denotes the target; magenta indicates ${\boldsymbol{\mathbf{u}}}(s)$ during ODE integration (Eq. 13); and blue shows the smooth trajectory ${\boldsymbol{\mathbf{z}}}(s)$ of the aerial manipulator.
5.2 Benchmark
We compare our method against two widely used sampling-based algorithms: RRT* (karaman2011sampling) and Model Predictive Path Integral control (MPPI) (williams2017information), as well as an ablation study that replaces SQs with ellipses for obstacle modeling. The comparisons are illustrated in Fig. 6 and Table 1. In the implementation, RRT* is sampled in ${\boldsymbol{\mathbf{z}}}_{sys}$, with collision detection based on the geometric overlap between the aerial manipulator and obstacles at each configuration $\boldsymbol{\mathbf{z}}$}. MPPI is evaluated with a cost function that penalizes both target distance and trajectory smoothness, using N = 50 rollouts per control step, which provides a trade-off between computational cost and trajectory quality. The ablation study adopts the same settings as the proposed algorithm, except that obstacles are represented as ellipses instead of SQs.
Figure 6: Comparison of end-effector trajectories by different planners: (a) the proposed SQ-based planner, (b) ablation study (ellipses) (c) RRT*, (d) MPPI.
Table 1: Benchmark with related planners
Method Time Min-distance Arc-length Jerkiness
Proposed 0.128 0.0644 1.375 0.0016
Ellipse 0.113 -0.0049 1.383 0.0001
RRT* 105.8 0.0314 1.565 0.0032
MPPI 0.552 0.0039 1.721 0.0057
We evaluate four performance metrics, as summarized in Table 1: i) Time denotes the online computational cost. For our method, this is the time to integrate Eq. 13. For MPPI, it corresponds to the time required for one receding-horizon planning cycle (i.e., one control update step). For RRT*, it is the average time needed to generate a collision-free path. ii) Minimum distance represents the closest distance between the robot system and obstacles. Our SQ-based method achieves maximum clearance, thus yielding the safest trajectories. In contrast, the ellipse-based ablation tends to overestimate obstacle sizes, which can trigger unnecessary avoidance maneuvers. This difference is particularly evident in cluttered environments. iii) Arc-length measures the total trajectory length. Shorter arc-length indicates a more efficient path. iv) Jerkiness quantifies trajectory smoothness. A lower value indicates a smoother motion, which is critical for stability.
6 Experimental Results
We also validate our planner and controller with an actual object-picking experiment using our aerial manipulator. Fig. 7 displays our experimental setting.
Figure 7: Flight experiment setup: (a) 3D side view in the laboratory environment. An industrial fan is utilized to simulate wind conditions. (b) 2D top view illustration.
6.1 Experimental Setups
The aerial manipulation platform combines a fully actuated hexacopter and a 3-DOF robot arm. The 3.50 kg hexacopter, based on a DJI F550 frame, uses six 9-inch APC LPB09045MR propellers, KDE2314XF-965 motors with KDEXF-UAS35 ESCs, and custom 3D-printed thrust-tilting mounts set at $\alpha_{p}$ = $15{}^{\circ}$. Power is supplied by separate 4S (Intel NUC) and 6S (ESCs) LiPo batteries. The arm uses three ROBOTIS Dynamixel XM430 actuators via a U2D2 interface. An Intel NUC running Ubuntu 20.04 and ROS Noetic handles control and OptiTrack-based navigation, while motor speeds are governed by a custom Pixhawk 4. An industrial fan is utilized to simulate wind conditions, and the parameter values are listed in Table 2.
Table 2: Parameters for planner and controller
$\alpha_{p}$, $L$, $k_{f}$ $\tfrac{\pi}{12}$ [rad], $0.278$ m, $0.016$ m
$\alpha$, $T_{d}$, $d_{0}$ 20 [-], $30$ s, $1.0$ mm
$(k_{\textrm{min}},k_{\textrm{max}})$ $(1.0\times 10^{-7},1.0\times 10^{3})$ N/m
[…]
$\boldsymbol{a_{0}}$, $\boldsymbol{a_{1}}$, $\boldsymbol{\varepsilon_{dob}}$ $\boldsymbol{I_{6}}$, $2\boldsymbol{I_{6}}$, $0.95\boldsymbol{I_{6}}$ [-]
$\alpha_{co}$, $\sigma_{co}$, $\underline{T}$, $\bar{T}$ 5.0 [-], 1.0 [-], 1.0 N, 15 N
$\boldsymbol{Q_{q}}$, $\boldsymbol{Q_{\ddot{\theta}}}$, $\boldsymbol{{\Gamma}_{q}}$, $\boldsymbol{\Gamma_{\theta}}$ $\textrm{blkdiag}\{\boldsymbol{I_{3}},3\boldsymbol{I_{3}}\}$, $4\boldsymbol{I_{3}}$, $4\boldsymbol{I_{6}}$, $5\boldsymbol{I_{3}}$ [-]
[…]
As illustrated in Fig. 7(a), obstacle 1 is a blue cuboid (0.4 m $\times$ 0.31 m $\times$ 1.03 m), obstacle 2 is a brown cuboid (0.48 m $\times$ 0.39 m $\times$ 1.27 m), and obstacle 3 is a white cylinder (radius 0.29 m, height 0.72 m). The poses of all obstacles and the target object are measured using the OptiTrack motion capture system.
6.2 Scenario
The aerial manipulator takes off and deploys the robot arm to its fully extended configuration ($\boldsymbol{\theta}=\boldsymbol{0_{3\times 1}}$). Subsequently, the path planning algorithm is activated, after which the vehicle follows the generated trajectory, guides the end-effector through the ring affixed to the target, ascends to grasp it, and returns to the initial position.
6.3 Result and Discussion
Figure 8: Histories of the multirotor’s measured and target displacements from initial position ($\Delta q_{1};\Delta q_{2};\Delta q_{3}]\triangleq[q_{1}-q_{1}(t_{0});q_{2}-q_{2}(t_{0});q_{3}-q_{3}(t_{0})]$ and $\Delta q_{t,1};\Delta q_{t,2};\Delta q_{t,3}]\triangleq[q_{t,1}-q_{t,1}(t_{0});q_{t,2}-q_{t,2}(t_{0});q_{t,3}-q_{t,3}(t_{0})]$), multirotor’s measured and target Euler angles ($q_{4};q_{5};q_{6}]$ and $q_{t,4};q_{t,5};q_{t,6}]$), robot arm’s measured and target joint angles ($\boldsymbol{\theta}$ and $\boldsymbol{\theta_{t}}$), collision avoidance CBF, $h_{co}$, and thrust values during approaching, picking, and returning.
Figs. 8 and 9 indicate that the task is completed without collisions. In Fig. 8, $h_{co}$ remains positive throughout the operation, confirming collision-free performance despite tracking errors in the multirotor’s position and attitude, and all motor thrusts adhere to their limits. These results verify the effectiveness of our controller for instantaneous collision avoidance. Fig. 9 displays the multirotor navigating and rotating within a narrow path, while the robot arm simultaneously bends to reach the target object. The attached video (https://youtu.be/hQYKwrWf1Ak) presents not only a more detailed visualization of the experimental process, but also illustrates how the CBF intervenes to block the target-approaching behavior.
Figure 9: 2D top view trajectories of the multirotor and the end-effector from the initial to picking poses.
7 Conclusion
This paper proposes whole-body motion planning and safety-critical control for an aerial manipulator using superquadrics and proxies, where both the multirotor platform and the robot arm linkages independently avoid collisions with surrounding obstacles. First, we derive the dynamic model of the aerial manipulator by attaching a multi-DOF robot arm to a fully actuated hexarotor. Second, we design a maximum-clearance whole-body motion planner by integrating the SQ–proxy representation with Voronoi diagrams on an equilibrium manifold. Third, a controller is developed to ensure thrust limits and collision avoidance. Through comparative simulations, we show that our framework outperforms sampling-based planners in cluttered environments, achieving faster, safer, and smoother trajectories, and surpasses ellipsoid-based ablation methods in accurately representing geometry. Additionally, we validate the dynamic feasibility and robustness of our approach on a physical aerial manipulator.
For future work, both the planning and control methods will be extended to three-dimensional environments and experimentally validated in more cluttered scenarios that include thin cylindrical obstacles only allowing narrow gaps near the target location.